OpenStudy (anonymous):

3^2 + (7-2) *4 - 6/3.... help?

OpenStudy (anonymous):

Follow the order of operations, starting with P for parenthesis.

OpenStudy (anonymous):

3^2 + (5) *4 - 6/3

OpenStudy (anonymous):

Now E for exponents

OpenStudy (anonymous):

9 + 5 *4 -6/3

OpenStudy (anonymous):

Good now M for multiply

HanAkoSolo (jamierox4ev3r):

just follow P.E.M.D.A.S, from left to right. (parentheses, exponents/roots, multiply, divide, addition, and subtraction.

OpenStudy (anonymous):

9 +20 -6/3

OpenStudy (anonymous):

Nice! now D for division

OpenStudy (anonymous):

9 + 20-2

HanAkoSolo (jamierox4ev3r):

from there it is just simple arithmetic.

OpenStudy (anonymous):

Ok! Now finish it with addition and subtraction

OpenStudy (anonymous):

27?

OpenStudy (anonymous):

Great job!

OpenStudy (anonymous):

thxs! :) ok! this ones kinda hard!

OpenStudy (anonymous):

[(21+6) -3^2]/9*2

OpenStudy (anonymous):

Just to clarify\[\frac{[\left( 21+6 \right)-3^{2}] }{ 9*2 }\]

OpenStudy (anonymous):

Is this right @ashleyjade ?

OpenStudy (anonymous):

umm yes!

OpenStudy (anonymous):

In that case, start with P again.

OpenStudy (anonymous):

(27-9)/9*2

OpenStudy (anonymous):

yep yep now P again

OpenStudy (anonymous):

17/9*2

OpenStudy (anonymous):

Right now M for multiply

OpenStudy (anonymous):

17/18?

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

i dont have an answer like that though

OpenStudy (anonymous):

a.) 4 b.) 4.6 c.)1 D.)52

OpenStudy (anonymous):

Oh I see 27-9=18 not 17

OpenStudy (anonymous):

so the answer is 1?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

one more? haha

OpenStudy (anonymous):

fire away

OpenStudy (anonymous):

ryans age can be described as 5 less then 2 times his cousins age. If his cousins age is represented by the variable "M", write the algrebraic expression represnting ryans age using m

OpenStudy (anonymous):

(2 times M) minus 5 = ryans age m 2M-5 = m

OpenStudy (anonymous):

thxs! :) whats ur name?

OpenStudy (anonymous):

Derek. Nice to meet you.

OpenStudy (anonymous):

nice meeting u! :)

OpenStudy (anonymous):

Keep at it Miss Jade. Then you can help others too.